### v2

# On second order conditions for singular optimal control of port-Hamiltonian systems

Abstract
We study singular optimal control problems of port-Hamiltonian (descriptor) systems. We employ general control-affine cost functionals that include as a special case the energy supplied to the system. We first derive optimality conditions for the case of ordinary differential equations with and without control bounds by applying the general theory to the specially structured port-Hamiltonian case, and show that this leads to nice optimality conditions, in particular in the linear case. We then extend these results to classes of port-Hamiltonian descriptor systems.
Keywords: singular optimal control, port-Hamiltonian descriptor system, Goh condition, Legendre-Clebsch condition, minimal supplied energy.
AMS subject classification: 49J15, 37J99, 34A09
1 Introduction
The energy based modeling of physical systems employing the model class of port-Hamiltonian descriptors, see e.g. [29, 43], has become an accepted modeling paradigm which has been successfully used in a multitude of applications from a wide variety of application domains: mechanics [5, 6, 24, 42], electrical engineering [37, 41], thermodynamics and fluid dynamics [2, 13, 35, 36], economics [23]. See [9, 14, 43] and the recent survey [29].
In this paper we consider a simplified form of real nonlinear port-Hamiltonian (pH) descriptor systems, see [28, 29], that have a governing equation
$E\dot{x}=\big{(}J(x)-R(x)\big{)}Qx+G(x)u(t),$ (1.1)
with the state $x$ belonging to the state space $\mathcal{X}:=W^{1,1}([0,T];\mathbb{R}^{n})$ (the Sobolev space of functions in $L^{1}[0,T]$ with weak first derivatives that have a finite $L^{1}$-norm) and the input (control) $u$ ranging in the control space $\mathcal{U}:=L^{1}([0,T];\mathbb{R}^{m})$ with $L^{1}$ denoting the usual Lebesgue space. Equation (1.1) has an associated energy function
$\mathcal{E}:\mathbb{R}^{n}\to\mathbb{R},\quad\ x\mapsto\frac{1}{2}x^{T}E^{T}Qx$ (1.2)
(often called Hamiltonian or storage function) that satisfies
$\frac{\partial\mathcal{E}}{\partial x}(x)=E^{T}Qx.$ (1.3)
The output $y$ is given by
$y=G^{T}(x)Qx.$ (1.4)
As is common in pH (descriptor) systems, the structure of the coefficients is the following:
  • $J:\mathbb{R}^{n}\to\mathbb{R}^{n,n}$ is skew-symmetric;
  • $R:\mathbb{R}^{n}\to\mathbb{R}^{n,n}$ is symmetric and positive semidefinite;
  • $E,Q\in\mathbb{R}^{n,n}$ are such that $E^{T}Q$ is symmetric and positive semidefinite and represents the Hessian of $\mathcal{E}=\frac{1}{2}x^{T}E^{T}Qx$;
  • $G:\mathbb{R}^{n}\to\mathbb{R}^{n,m}$.
Note that if $E$ is invertible, then one obtains the classical form of a pH system by a change of variables (see [43]).
The success of modeling within the class of pH (descriptor) systems is due to its many important properties which include the invariance of the class under power-conserving interconnection, which greatly simplifies modularized automated modeling, and the invariance under Galerkin projection which makes the class very suitable for discretization and model reduction. Physical properties like energy dissipation, Lyapunov stability and passivity are encoded in the structure of the equations leading directly to another key property of port-Hamiltonian (descriptor) systems, the power balance equation and the resulting dissipation inequality, see e.g. [28, 29] or [43] for the case of ordinary pH systems. From a physics point of view, pH systems model the interaction of three types of energy/power. The stored energy (in the energy storing components) is presented by the nonnegative quadratic form $\mathcal{E}(x)$, the dissipated energy by the nonnegative quadratic form $\mathcal{D}(x)=-x^{T}Q^{T}RQx$ and the supplied energy by the form $\mathcal{S}(y,u)=y^{T}u$.
Example 1.1.
A classical example of pH modeling are (dissipative) Hamiltonian equations of motion, see e.g. [1, 43], which, in first order descriptor representation using the position coordinate $q$ and velocity coordinate $\tilde{p}$ (replacing the usual linear momentum $p=M\dot{q}$), together with a force term $\tilde{B}(q)u$, take the descriptor form
$\begin{bmatrix}M&0\\ 0&I\end{bmatrix}\begin{bmatrix}\dot{\tilde{p}}\\ \dot{q}\end{bmatrix}=\begin{bmatrix}-D(p,q)&-I\\ I&0\end{bmatrix}\begin{bmatrix}I&0\\ 0&K\end{bmatrix}\begin{bmatrix}\tilde{p}\\ q\end{bmatrix}+\begin{bmatrix}\tilde{B}(q)\\ 0\end{bmatrix}u$ (1.5)
coupled with a collocated output
$y=\begin{bmatrix}\tilde{B}(q)^{T}&0\end{bmatrix}\begin{bmatrix}I&0\\ 0&K\end{bmatrix}\begin{bmatrix}\tilde{p}\\ q\end{bmatrix}.$ (1.6)
Here the energy function (Hamiltonian) is
$\mathcal{E}(q,\tilde{p})=\frac{1}{2}{\tilde{p}}^{T}M\tilde{p}+\frac{1}{2}q^{T}Kq,$
i.e., kinetic plus potential energy, where $M$ is a positive (semi-)definite mass matrix, $K$ is a positive definite stiffness matrix, and $D(q,p)=D(q,p)^{T}$ is a positive semidefinite coefficient matrix that models internal damping. This representation has the advantage that small masses can be set to 00, leading to a singular mass matrix $M$. See [30] for a detailed analysis of different representations and transformations between the different formulations. A typical optimal control application is arising in the control of the system into an equilibrium state, e.g. in the stabilization of buildings excited by earthquakes, see e.g. [25].
With an increased understanding of the many advantageous properties of modeling with pH (descriptor) systems, in recent years also feedback control and optimal control problems for this model class have become an important research topic in different application fields [8, 10, 11, 17, 19, 26, 31, 32, 38, 39, 45].
In the optimal control of pH (descriptor) systems, a natural cost function is the integral of the supplied energy
[…]
It is readily available as a mathematical expression in terms of the ports and it is interesting from an application point of view to employ it as the objective to be minimized in optimal control problems (OCPs) [10, 11, 32, 39]. However, using the supplied energy as cost function, the resulting OCP is typically singular, it is missing the (commonly used) positive quadratic regularization term in $u$, and standard solution techniques, for instance the construction of a Riccati state feedback, see e.g. [21, 27], are not directly applicable. In view of this difficulty, the topic of this paper is the analysis of singular OCPs with constraints given by pH (descriptor) systems without the regularization approach.
We obtain our results by applying optimality conditions for singular optimal control of control-affine systems. More precisely, considering problems of the form
$\min$ $\quad\Psi(x(T))$ (1.7)
$\mathrm{s.t.}$ $\quad\dot{x}=f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i},$ (1.8)
$\quad x(0)=x_{0}.$ (1.9)
with $f_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, for $i=0\ldots,m$ and $\Psi:\mathbb{R}^{n}\to\mathbb{R},$ one gets the well-known Goh conditions which state that, along optimal trajectories,
$p^{T}[f_{i},f_{j}]=0,\quad\text{for }i,j=1,\dots,m,$ (1.10)
where $p$ is the adjoint state, and where, for any pair of smooth vector fields $h,k\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$, $[h,k]:=k^{\prime}h-h^{\prime}k$ is their Lie bracket, with $k^{\prime},h^{\prime}$ denoting the Jacobians of $k,h$. Moreover, the generalized Legendre-Clebsch conditions ensure that the matrix with entries
$p^{T}[f_{j},[f_{0},f_{i}]]$ (1.11)
is negative semidefinite and, whenever it is definite, it provides a feedback formula for the singular control. These results are detailed in Appendix A.
At this point we recall that any problem with a general control-affine integral cost can be rewritten in the Mayer form (1.7)-(1.9) by adding an extra scalar state variable (see (2.1) below). Therefore, the formulation (1.7)-(1.9) covers problems with integral-type costs of the form $\mathcal{J}\coloneqq\int_{0}^{T}\ell_{0}(x)+\sum_{i=1}^{m}\ell_{i}(x)u_{i}\,dt,$ where $\ell_{i}:\mathbb{R}^{n}\to\mathbb{R}$ for $i=0\ldots,m$.
The paper is organized as follows.
In Appendix A we recall second order optimality conditions for problems of the form (1.7)-(1.9). In Section 2 we analyze how these general optimality conditions look when we specialize the constraint function from general ordinary differential equations to ordinary port-Hamiltonian systems and we show that the structure helps to obtain much more elegant conditions. We extend these results to port-Hamiltonian descriptor systems in Section 3 and we present some conclusions and directions of future work in Section 4.
2 Optimality conditions for nonlinear pH descriptor systems
The optimality conditions for singular optimal control problems that are presented in Appendix A do not exploit the pH structure. In this section, we therefore make explicit use of the nonlinear pH structure in (1.1) and generalize results of [10, 26, 39].
2.1 Optimality conditions for ordinary nonlinear pH systems
We first consider the following optimal control problem governed by the nonlinear ordinary pH dynamics (1.1) with $E=I$, minimizing a general control-affine integral cost
$\min$ $x_{n+1}(T)$ (2.1)
$\mathrm{s.t.}$ $\dot{x}=\big{(}J(x)-R(x)\big{)}\frac{\partial\mathcal{E}}{\partial x}(x)+G(x)u(t),$
$\dot{x}_{n+1}=\ell_{0}(x)+\sum_{i=1}^{m}\ell_{i}(x)u_{i},$
$\begin{bmatrix}x\\ x_{n+1}\end{bmatrix}(t_{0})=\begin{bmatrix}x_{0}\\ 0\end{bmatrix},$
with the assumption that the control $u$ takes values in an open set $\mathbb{U}\subseteq\mathbb{R}^{m}$. For the sake of simplicity in the presentation of the results, we introduce the abbreviations
$g_{0}(x)\coloneqq\big{(}J(x)-R(x)\big{)}Qx=\big{(}J(x)-R(x)\big{)}\frac{\partial{\mathcal{E}}}{\partial x}(x),$ (2.2)
and, setting $z\coloneqq\begin{bmatrix}x\\ x_{n+1}\end{bmatrix}$ and writing $G=\begin{bmatrix}g_{1}\,\dots\,g_{m}\end{bmatrix}$, we define
$f_{0}(z)\coloneqq\begin{bmatrix}g_{0}(x)\\ \ell_{0}(x)\end{bmatrix};\qquad f_{i}(z)\coloneqq\begin{bmatrix}g_{i}(x)\\ \ell_{i}(x)\end{bmatrix}.$ (2.3)
In this case, recalling the Hamilton function introduced in (A.1), the adjoint variable has an additional coordinate $p_{n+1}$ and takes the form $\begin{bmatrix}p\\ p_{n+1}\end{bmatrix},$ where $p$ maps into $\mathbb{R}^{n}$ and $p_{n+1}$ is scalar. In the absence of (state or terminal) constraints, the multiplier $p_{n+1}$ is unique and the associated extremal is normal, i.e., the corresponding multiplier $p_{n+1}$ can be set to $1$. Therefore, throughout the remainder of the article, we fix $p_{n+1}=1$.
Theorem 2.1 (Goh conditions for nonlinear ordinary pH control).
For the optimal control problem with nonlinear pH constraint (2.1), it holds that
$\left\{\begin{array}[]{rl}p^{T}[g_{i},g_{j}]&=0,\\ \ell_{j}^{\prime}\,g_{i}-\ell_{i}^{\prime}\,g_{j}&=0,\end{array}\right.\qquad\text{for }\,i,j=1,\ldots,m,$
along optimal trajectories.
Proof.
The result follows from Theorem A.1 Appendix A. ∎
Using latter result, we compute the expressions for the generalized Legendre-Clebsch conditions (see Theorem A.2 in Appendix A, and $W$ and $d$ given in (A.8)) in the context of problem (1.7)–(1.9). We have
$[f_{0},f_{i}]=\begin{bmatrix}[g_{0},g_{i}]\\ \ell_{i}^{\prime}g_{0}-\ell_{0}^{\prime}g_{i}\end{bmatrix},$ (2.4)
and
$[f_{j},[f_{0},f_{i}]]=\\ \begin{bmatrix}[g_{j},[g_{0},g_{i}]]\\ \ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{j}]\end{bmatrix}.$
We set
$w_{ij}\coloneqq p^{T}[g_{j},[g_{0},g_{i}]]\\ +\Big{(}\ell_{i}^{\prime\prime}g_{0}g_{j}+\ell_{i}^{\prime}g_{0}^{\prime}g_{j}-\ell_{0}^{\prime\prime}g_{i}g_{j}-\ell_{0}^{\prime}g_{i}^{\prime}g_{j}-\ell_{j}^{\prime}[g_{0},g_{i}]\Big{)},$ (2.5)
and
$d_{i}\coloneqq p^{T}[g_{0},[g_{0},g_{i}]]\\ +\Big{(}\ell_{i}^{\prime\prime}g_{0}g_{0}+\ell_{i}^{\prime}g_{0}^{\prime}g_{0}-\ell_{0}^{\prime\prime}g_{i}g_{0}-\ell_{0}^{\prime}g_{i}^{\prime}g_{0}-\ell_{0}^{\prime}[g_{0},g_{i}]\Big{)},$ (2.6)
and obtain the following result.
Theorem 2.2 (Legendre-Clebsch condition for nonlinear ordinary pH control).
For the nonlinear pH optimal control problem (2.1), the matrix $W=[w_{ij}],$ with $w_{ij}$ given in (2.5), is negative semidefinite. Additionally, whenever $W$ is negative definite over some control interval then, on that interval, one can write the optimal control in feedback form via the expression
$Wu+d=0.$ (2.7)
Proof.
The proof follows from Theorem A.2 in Appendix A. ∎
2.2 Optimality conditions for nonlinear ordinary pH systems with control bounds
In this section we discuss problem (2.1) in the presence of bounds on the control. This means that we no longer assume that $\mathbb{U}$ is open, but consider a set of admissible control values
$\mathbb{U}=[\underline{u}_{1},\overline{u}_{1}]\times\dots[\underline{u}_{m},\overline{u}_{m}],$ (2.8)
with real numbers $\underline{u}_{i}\leq\overline{u}_{i},$ for each $i=1,\dots,m$. For the problem with control bounds in the port-Hamiltonian optimal control problem (2.1), the Goh conditions of Theorem 2.1 specialize as described next.
Theorem 2.3 (Goh condition for ordinary pH control under control constraints).
Consider the nonlinear pH optimal control problem (2.1) under the control constraint $u(t)\in\mathbb{U}$, for a.e. $t\in[0,T],$ with the set $\mathbb{U}$ as in (2.8). Then, on any open interval over which an optimal control $u^{*}$ verifies
$\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i}\quad\text{ and }\quad\underline{u}_{j}<u^{*}_{j}(t)<\overline{u}_{j},$
for some pair $i,j=1,\ldots,m,$ it must necessary hold that
$p^{T}[g_{i},g_{j}]=0\quad\text{ and }\quad\ell_{j}^{\prime}\,g_{i}-\ell_{i}^{\prime}\,g_{j}=0.$
In view of (2.4)-(2.5), the results in Theorem 2.2 with control constraints take the following form.
Theorem 2.4 (Generalized Legendre-Clebsch condition and feedback formula for nonlinear ordinary pH control under control constraints).
Under the settings of Theorem 2.3, one gets that
$p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j\in S}u_{j}^{*}p^{T}[f_{j},[f_{0},f_{i}]]$
$+\sum_{j\in\underline{B}}\underline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]+\sum_{j\in\overline{B}}\overline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]=0.$
holds on singular arcs, where the expression of $w_{ij}\coloneqq p^{T}[f_{0},[f_{0},f_{i}]]$ is given by (2.5) above, $i\in S$ and $S$, $\underline{B}$ and $\overline{B}$ denote the set of singular, lower-bang and upper-bang component of $u^{*}$ In other words, Theorem A.4 in Appendix A holds.
Proof.
See Theorem A.4 in Appendix A. ∎
Example 2.5.
In Example 1.1, with the associated cost of minimizing the energy supply
[…]
the Lie brackets $[f_{i},f_{j}]$ are given by
$[f_{i},f_{j}]=\begin{bmatrix}0\\ 0\\ B_{j}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{i}(q)-B_{i}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{j}(q)\end{bmatrix}.$
These vanish identically for any sufficiently regular $\mathcal{E}$. In the current framework, one has $\mathcal{E}{\partial p^{2}}=M^{-1},$ for $M$ being symmetric positive definite, so Goh conditions (1.10) hold trivially. One gets the following expression for the second-order Lie brackets
$[f_{j},[f_{0},f_{i}]]=\begin{bmatrix}B_{i}^{\prime}M^{-1}B_{j}+B_{j}^{\prime}M^{-1}B_{i}\\ 0\end{bmatrix}\\ +\begin{bmatrix}\frac{\partial D}{\partial p}(M^{-1}B_{i})B_{j}+\frac{\partial^{2}D}{\partial p^{2}}(M^{-1}p,B_{i})B_{j}\\ 0\end{bmatrix},$
which considerably simplifies under additional assumptions on $D$ or $B$, and may lead to very simple expressions for the singular controls.
2.3 Optimality conditions for nonlinear pH descriptor systems
In this section we extend the results from the previous two subsections to a special class of pH descriptor systems. For general descriptor systems optimality condition have been derived in [20], see [22] for a detailed exposition. But again these do not exploit the particular structure.
It has been shown in [7] that by an appropriate regularization procedure, using transformations, derivatives, and feedback designs, a general descriptor system can always be transformed to a so-called index-one or strangeness-free system, see [22].
Then for a system of the form (1.1)–(1.4), after appropriate coordinate changes, we may assume that the system has the structure
$\begin{bmatrix}I_{n_{1}}&0\\ 0&0\end{bmatrix}\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}G_{1}\\ G_{2}\end{bmatrix}(x)\,u+$ (2.9)
$\qquad\left(\begin{bmatrix}J_{11}&J_{12}\\ -J_{12}^{T}&J_{22}\end{bmatrix}(x)-\begin{bmatrix}R_{11}&R_{12}\\ R_{12}^{T}&R_{22}\end{bmatrix}(x)\right)\begin{bmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$
with output
$y=\begin{bmatrix}G_{1}(x)^{\top}&G_{2}(x)^{\top}\end{bmatrix}\begin{bmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}.$ (2.10)
Observe that the condition $Q^{T}E=E^{T}Q\geq 0$ implies that $Q_{12}=Q_{21}^{T}=0$ and $Q_{11}\geq 0$. In the following, to keep the presentation simple, we also assume that $G_{2}(x)=0$. If this is not the case then the following reduction procedure will lead to a system with a feed-through term for which an extra transformation is necessary, see [29]. The assumption that the system is of index at most one means that from the algebraic equation
$0=$ $[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{21}$
$+(-J_{12}^{T}(x_{1},x_{2})-R_{12}^{T}(x_{1},x_{2}))Q_{11}x_{1}+[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{22}x_{2},$
the variable $x_{2}$ can be expressed via the Implicit Function Theorem as a function of $x_{1}$ and inserted in the coefficients. Due to the structure of $G(x)$, $x_{2}$ contributes to the cost functional only indirectly as argument of $G$. Thus, we can first solve the optimal control problem for $x_{1}$ and obtain $x_{2}$ in a post-processing step. This means that the dynamics of the system is described by the ordinary pH system
$\dot{x}_{1}$ $=(\tilde{J}_{11}(x_{1})-\tilde{R}_{11}(x_{1}))Q_{11}x_{1}+\tilde{G}_{1}(x_{1})u,$
$y$ $=\tilde{G}_{1}^{T}(x_{1})Q_{11}x_{1},$
and the cost functional is given by
[…]
Thus we have reduced the problem to one of the form (2.1) for which the previous simplified formulas directly apply.
3 Optimality conditions for linear pH descriptor systems
In the last section we have presented the Goh conditions for pH (descriptor) systems which turned out to be much simpler than those for general systems. This simplicity is even more pronounced in the case of linear pH systems. We again start with the optimal control of ordinary linear pH systems.
3.1 Optimality conditions for linear pH systems without control constraints
In this subsection we analyze the Goh optimality conditions for problems when the constraint equation is given by a linear time-invariant pH system and the energy function $\mathcal{E}=\frac{1}{2}x^{T}Qx$ is again quadratic in the state, i.e.,
$\dot{x}(t)$ $=(J-R)Qx(t)+Gu(t),\quad x(0)=x^{0},$ (3.1)
$y(t)$ $=G^{T}Qx(t).$
where
  • $J\in\mathbb{R}^{n,n}$ is skew-symmetric,
  • $R\in\mathbb{R}^{n,n}$ is symmetric positive semidefinite,
  • $Q\in\mathbb{R}^{n,n}$ is symmetric positive definite,
  • $G=\begin{bmatrix}g_{1}&\ldots&g_{m}\end{bmatrix}\in\mathbb{R}^{n,m}$ has full rank $m\leq n,$
  • $u(t)\in\mathbb{U}$ for a.e. $t,$ where $\mathbb{U}$ is an open subset of $\mathbb{R}^{m}.$
Note that more general linear port-Hamiltonian formulations are possible, we follow here the formulation considered in [10, 11, 39].
We study the optimal control problem with quadratic cost
[…]
where the matrices involved in the cost functional satisfy
  • $Y\in\mathbb{R}^{n,n}$ is symmetric and positive semidefinite,
  • $N=\begin{bmatrix}n_{1}&\ldots&n_{m}\end{bmatrix}\in\mathbb{R}^{m,m}$,
  • $\ell=\begin{bmatrix}\ell_{1}&\ldots&\ell_{m}\end{bmatrix}^{T}\in\mathbb{R}^{m}$.
In the following, for simplicity, we often use the abbreviation
$A\coloneqq(J-R)Q.$ (3.3)
To compute the Lie bracket expressions for problem (3.2), we again add an auxiliary state variable $x_{n+1}$ to transform the problem into Mayer form (with only terminal cost) given by the following dynamics
$\dot{x}_{n+1}$ $=y^{T}Yy+y^{T}Nu+\ell^{T}u=x^{T}QGYG^{T}Qx+x^{T}QGNu+\ell^{T}u,$
$x_{n+1}(0)$ $=0.$
Then we set
$f_{0}:=\begin{bmatrix}Ax\\ x^{T}QGYG^{T}Qx\end{bmatrix},\quad f_{i}:=\begin{bmatrix}g_{i}\\ x^{T}QGn_{i}+\ell_{i}\end{bmatrix},$
for $i=1,\dots,m,$ thus $f_{0},f_{1},\dots,f_{m}$ are functions from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$. We obtain
$[f_{i},f_{j}]=\begin{bmatrix}0\\ \vdots\\ 0\\ n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qg_{j}\end{bmatrix}\quad i,j=1,\dots,m,$
so that
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j}\quad i,j=1,\dots,m,$ (3.4)
and we get the following Goh condition.
Theorem 3.1 (Goh condition for linear pH systems).
For the optimal control problem (3.2),
$N^{T}G^{T}QG\text{ is symmetric}$ (3.5)
along optimal trajectories.
Proof.
The result is a direct consequence of Theorem A.1 in the Appendix A. ∎
An optimality condition for linear pH systems has been studied in [39]. In order to compare our results with those in [39], we see that the optimality condition (3.5) is automatically satisfied in their framework, since there $N=I_{m}$ is a standing hypothesis. Additionally, observe that we do not need the normality hypothesis used in [39], since our problem (3.2) does not involve terminal constraints, while in [39] the terminal point is fixed.
We next compute the expressions of the Lie brackets for the linear problem (3.2) in order to display the Goh and Legendre-Clebsch conditions. For $i=1,\dots,m$ we get
$[f_{0},f_{i}]=\begin{bmatrix}-Ag_{i}\\ n_{i}^{T}G^{T}QAx-x^{T}QG(M+M^{T})G^{T}Qg_{i}\end{bmatrix},$
$\left[f_{0},[f_{0},f_{i}]\right]=\begin{bmatrix}A^{2}g_{i}\\ n_{i}^{T}G^{T}QA^{2}x-2g_{i}^{T}QGMG^{T}QAx+2x^{T}QGMG^{T}QAg_{i}\end{bmatrix},$
and
$d_{i}$ $=$ $\begin{bmatrix}p\\ 1\end{bmatrix}^{T}\left[f_{0},[f_{0},f_{i}]\right]$
$=$ $p^{T}A^{2}g_{i}+n_{i}^{T}G^{T}QA^{2}x-2g_{i}^{T}QGMG^{T}QAx+2x^{T}QGMG^{T}QAg_{i}.$
Additionally, for $i,j=1,\dots,m$, we have
$\left[f_{j},[f_{0},f_{i}]\right]=\begin{bmatrix}0\\ n_{i}^{T}G^{T}QAg_{j}-2g_{i}^{T}QGMG^{T}Qb_{j}+n_{j}^{T}G^{T}QAg_{i}\end{bmatrix}$
and
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}\left[f_{j},[f_{0},f_{i}]\right]=n_{i}^{T}G^{T}QAG_{j}-2g_{i}^{T}QGMG^{T}Qg_{j}+n_{j}^{T}G^{T}QAg_{i}.$
Consequently, the matrix $W$ in (3.2) is given by
$W=N^{T}G^{T}QAG-2G^{T}QGMG^{T}QG+(N^{T}G^{T}QAG)^{T}.$
Inserting the structure of the matrix $A$ (see (3.3)) and using the skew-symmetry of $J$, we get
$W=-N^{T}G^{T}QRQG-2G^{T}QGMG^{T}QG-(N^{T}G^{T}QRQG)^{T},$ (3.6)
and we obtain the following Legendre-Clebsch condition.
Theorem 3.2 (Generalized Legendre-Clebsch condition for linear pH problems).
Whenever the optimal control problem (3.2) of the linear port-Hamiltonian admits an optimal solution, the matrix $W$ given in (3.6) is negative semidefinite.
In the often considered special case of minimizing the supplied energy, that is, setting
$M=0,\quad\ell=0,\quad N=I,$ (3.7)
we then get
$W=-2G^{T}QRQG,$ (3.8)
which is positive semidefinite in view of the standing hypotheses for $Q$ and $R$. Moreover, whenever $R$ is definite, so will be $W,$ leading then to a feedback expression for the singular control from (2.7). In the latter case, along such trajectories, the solution is asymptotically stable since the dissipation term acts in this direction. Additionally, under the hypotheses used in [39], namely, (3.7) together with
${\rm Im}(G)\cap{\rm Ker}(RQ)=0,$ (3.9)
the expression of $W$ also coincides with (3.8), which is negative definite in view of assumption (3.9). Here we see directly that under the conditions (3.7) and (3.9) (as in [39]), a feedback solution exists. On the other hand our analysis is more general, since we get analogous results under noticeably weaker assumptions.
Remark 3.1.
In Theorems 3.1 and 3.2, we have derived the Goh and generalized Legendre-Clebsch conditions for linear port-Hamiltonian (pH) systems, which correspond to second-order optimality conditions. Higher-order conditions (see, e.g., [16, 18]) involve increasingly complex higher-order Lie brackets. Even for linear pH systems, these expressions become overly intricate, limiting their practical utility and insight.
Remark 3.2 (Cost functional with reference trajectory).
In many applications one wants to control a system towards a given reference trajectory $x_{\rm ref}$. Replacing $x=\tilde{x}-x_{\rm ref}$ one gets a given inhomogeneity in the state and the output equation. In this case, the system results in
$\dot{\tilde{x}}$ $=(J-R)Q\tilde{x}+Gu+h(t),$ (3.10)
$y$ $=G^{T}Q(\tilde{x}-x_{\rm ref})=G^{T}Q\tilde{x}+k(t).$
After removing the terms depending only on the reference trajectory, the cost functional takes the form
$\int_{0}^{T}\big{[}$ $(x-2x_{\rm ref})^{T}QGYG^{T}Qx+(x-x_{\rm ref})^{T}QGNu+\ell^{T}u\big{]}\,dt.$ (3.11)
For the optimal control problem governed by system (3.10) and associated to the cost (3.11), the expressions for the Lie brackets coincides with (3.4), the matrix $W$ matches with (3.6), so that the optimality conditions do not change. Additionally, the vector $d$ is given by
$d_{i}=$ $p^{T}A^{2}g_{i}+n_{i}^{T}G^{T}QA^{2}x-2g_{i}^{T}QGYG^{T}QAx$
$+2(x-x_{\rm ref})^{T}QGYG^{T}QAg_{i}.$
Example 3.3.
To illustrate the elegance of the optimality conditions, consider the following example of a second order linear control problem that is discussed in detail in [10] and arises e.g. in the control of high rise buildings [44]:
$M\tfrac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}{q}+D\tfrac{\mathrm{d}}{\mathrm{d}t}{q}+Kq=Bu$ (3.12)
with real symmetric positive definite matrices $M,K$ and $B$ of full column rank. In contrast to [10] we assume that the damping does not work on the whole position vector, i.e., that $D$ is only positive semidefinite. Setting $x=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}=\begin{bmatrix}M\dot{q}\\ q\end{bmatrix}$, $R=\begin{bmatrix}D&0\\ 0&0\end{bmatrix}$, $J=\begin{bmatrix}0&-I\\ I&0\end{bmatrix}$, $Q=\begin{bmatrix}M^{-1}&0\\ 0&K\end{bmatrix}$, $G=\begin{bmatrix}B\\ 0\end{bmatrix}$, and introducing a collocated output $y=G^{T}Qx$, the system has the form (3.1). Unfortunately the condition ${\rm Im}(G)\cap{\rm Ker}(RQ)=\{0\}$ that is used in [39] does not hold, since $DM^{-1}$ has a kernel and $W=-2B^{T}M^{-1}DM^{-1}B$ is only semidefinite. Hence the techniques in  [39] cannot be applied. Also the linear system with $W$ for the feedback is, in general, not (uniquely) solvable. However, still a (non-unique) feedback solution may exist if the right hand side $d$ is in the image of $W$, see [10, 27]. In case $M$ is only semidefinite then one can rewrite the system in descriptor form as in (1.5) and the presented procedures apply to this case. We then obtain the optimality conditions as in Section 2.3.
In this subsection we have seen that by using the general theory we can generalize the conditions for optimality presented in [39]. In the next subsection we consider the case of control bounds.
3.2 Optimality conditions for linear pH problems subject to control bounds
If we require control bounds, i.e.
$\mathbb{U}=[\underline{u}_{1},\overline{u}_{1}]\times\dots[\underline{u}_{m},\overline{u}_{m}],$
so that $u$ is subject to the inequalities
$\underline{u}_{i}\leq u_{i}(t)\leq\bar{u}_{i},\quad i=1,\dots,m,$ (3.13)
then we obtain the optimal control problem
$\begin{split}\min\,\,&\int_{0}^{T}y^{T}Yy+y^{T}Nu+\ell^{T}u\\ {\rm s.t.}\,\,\,\,&\dot{x}(t)=(J-R)Qx(t)+Gu(t),\\ &x(0)=x^{0},\\ &y(t)=G^{T}Qx(t),\\ &\underline{u}_{i}\leq u_{i}(t)\leq\bar{u}_{i},\quad i=1,\dots,m.\end{split}$ (3.14)
In order to study (3.14), we consider switching functions. For the $i$-th component of the control, the switching function is given by
$s_{i}\coloneqq\mathcal{H}_{u_{i}}=\begin{bmatrix}p\\ 1\end{bmatrix}^{T}f_{i}=p^{T}g_{i}+x^{T}QGn_{i}+\ell_{i}.$ (3.15)
Introduce the set of switching points
$Z_{i}\coloneqq\{t\in[0,T]:s_{i}(t)=0\},$
and the interior set for control variable $i,$ for $i=1,\dots,m:$
$W_{i}\coloneqq\{t\in[0,T]:\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i}\}.$
In view of the maximum condition of Pontryagin’s Maximum Principle [33], one then has that $W_{i}$ is contained, up to a set of measure zero, in $Z_{i}$.
Now, following the notation in [39, Theorem 8], for any subset $\mathcal{I}\subseteq\{1,\dots,m\}$, we set
$u_{\mathcal{I}}\coloneqq(u_{i})^{T}_{i\in\mathcal{I}},\qquad G_{\mathcal{I}}\coloneqq(g_{i})^{T}_{i\in\mathcal{I}},\qquad s_{\mathcal{I}}\coloneqq(s_{i})^{T}_{i\in\mathcal{I}},$
and analogous notations are used for other involved matrices and vectors.
At this point, it is useful to present the adjoint equation for (3.14) which reads
$\left\{\begin{array}[]{rl}\displaystyle\frac{d}{dt}p^{T}&=-p^{T}A-p_{n+1}\left(x^{T}QG(Y+Y^{T})G^{T}Q+u^{T}N^{T}G^{T}Q\right),\\[4.30554pt] \dot{p}_{n+1}&=0,\end{array}\right.$
recalling that $p_{n+1}\equiv 1,$ since the involved terminal cost is $x_{n+1}(T).$
Theorem 3.4 (Goh and generalized Legendre-Clebsch conditions for linear pH problems under control constraints).
For problem (3.14), the Goh conditions read
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j}=0,$ (3.16)
on any open interval $(a,b)$ on which
$\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i}\quad\text{ and }\quad\underline{u}_{j}<u^{*}_{j}(t)<\overline{u}_{j}.$
The generalized Legendre-Clebsch condition states that the matrix
$N_{\mathcal{I}}^{T}G^{T}QAB_{\mathcal{I}}-2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}}+\left(N_{\mathcal{I}}^{T}G^{T}QAG_{\mathcal{I}}\right)^{T}$ (3.17)
is negative semidefinite on any open interval $(a,b)\subset[0,T]$ that is contained in $\bigcap_{i\in\mathcal{I}}W_{i}$ for $\mathcal{I}\subseteq\{1,\dots,m\}$, i.e., $\mathcal{I}$ is any subset of indexes of simultaneous singular components of the control.
Proof.
The proof follows from Theorems A.3 and A.4 in the Appendix. ∎
Example 3.5 (Special cases $N=0$ and $N=I_{m}$).
When $N=0$ or $N=I_{m}$, using the notation introduced in (2.3) for the vector fields $f_{i},$ one has that $[f_{i},f_{j}]=0$ in the whole space $\mathbb{R}^{n},$ independently of the trajectories. Therefore, the Goh conditions hold trivially. Additionally, for $N=0,$ the matrix in the generalized Legendre-Clebsch condition of Theorem 3.4 reads
$-2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}},$ (3.18)
and for $N=I_{m},$ it has the form
$-2G_{\mathcal{I}}^{T}Q[R+GYG^{T}]QG_{\mathcal{I}},$ (3.19)
We obtain the following feedback formulas.
Theorem 3.6 (Feedback formulas for singular controls).
For any subset $\mathcal{I}\subseteq\{1,\dots,m\},$ given an open interval $(a,b)\subseteq\bigcap_{i\in\mathcal{I}}W_{i},$ the following expressions for the control hold along optimal trajectories of problem (3.14)
$0=p^{T}A^{2}G_{\mathcal{I}}$
$+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
$+u^{T}_{\mathcal{I}}\left[N_{\mathcal{I}}^{T}G^{T}QAG_{\mathcal{I}}-2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}}+\left(N_{\mathcal{I}}^{T}G^{T}QAG_{\mathcal{I}}\right)^{T}\right]$
$+u^{T}_{\mathcal{A}}\left[N_{\mathcal{A}}^{T}G^{T}QAG_{\mathcal{I}}-2G_{\mathcal{A}}^{T}QGYG^{T}QG_{\mathcal{I}}+G_{\mathcal{A}}^{T}A^{T}QGN_{\mathcal{I}}\right].$
Additionally, for the special case $N=0$, one gets the reduced expression
00 0 $=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
$\quad-2(u^{T}_{\mathcal{I}}G_{\mathcal{I}}^{T}+u^{T}_{\mathcal{A}}G_{\mathcal{A}}^{T})QGYG^{T}QG_{\mathcal{I}},$
and for $N=I_{m}$,
00 0 $=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
$\quad-2(u^{T}_{\mathcal{I}}G_{\mathcal{I}}^{T}+u^{T}_{\mathcal{A}}G_{\mathcal{A}}^{T})Q\left[R+GYG^{T}\right]QG_{\mathcal{I}}.$
Proof.
Following the notation introduced above, and from (3.15), one gets
$s_{\mathcal{I}}=p^{T}G_{\mathcal{I}}+p_{n+1}\left(x^{T}QGN_{\mathcal{I}}+\ell_{\mathcal{I}}\right).$
Therefore,
$\dot{s}_{\mathcal{I}}$ $=$ $\frac{d}{dt}p^{T}G_{\mathcal{I}}+p_{n+1}\dot{x}^{T}QGN_{\mathcal{I}}$
$=$ $-p^{T}AG_{\mathcal{I}}-p_{n+1}x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}$
$-u^{T}\overbrace{N^{T}G^{T}QG_{\mathcal{I}}}^{(N^{T}G^{T}QG)_{\mathcal{I}}}+x^{T}A^{T}QGN_{\mathcal{I}}+p_{n+1}u^{T}\underbrace{G^{T}QGN_{\mathcal{I}}}_{(G^{T}QGN_{\mathcal{I})}}.$
We split the remainder of the proof in two cases.
Case 1: $N=0$ or $N=I_{m}$. From (3.2) one gets that
$\dot{s}_{\mathcal{I}}=-p^{T}AG_{\mathcal{I}}-x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}+x^{T}A^{T}QGN_{\mathcal{I}}.$ (3.21)
Thus,
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$ (3.22)
$+$ $u^{T}\left[N^{T}G^{T}QAG_{\mathcal{I}}-2G^{T}QGYG^{T}QG_{\mathcal{I}}+G^{T}A^{T}QGN_{\mathcal{I}}\right].$
For $N=0$, splitting the control vector in $u_{\mathcal{I}}$ and $u_{\mathcal{A}},$ one gets from latter equation that
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
$-2(u^{T}_{\mathcal{I}}G_{\mathcal{I}}^{T}+u^{T}_{\mathcal{A}}G_{\mathcal{A}}^{T})QGYG^{T}QG_{\mathcal{I}},$
where $\mathcal{A}\coloneqq\mathcal{I}^{c}.$
For $N=I_{m}$, from (3.22), one derives
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
-- $2(u^{T}_{\mathcal{I}}G_{\mathcal{I}}^{T}+u^{T}_{\mathcal{A}}G_{\mathcal{A}}^{T})Q\left[R+GYG^{T}\right]QG_{\mathcal{I}},$
where we replaced the matrix $A$ by its expression in (3.3).
Case 2: Arbitrary matrix $N$. From expression (3.2) for $\dot{s}_{\mathcal{I}}$, one obtains
$\ddot{s}_{\mathcal{I}}$ $=p^{T}A^{2}G_{\mathcal{I}}$
$+$ $x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
$+$ $u^{T}_{\mathcal{I}}\left[N_{\mathcal{I}}^{T}G^{T}QAG_{\mathcal{I}}-2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}}+\left(N_{\mathcal{I}}^{T}G^{T}QAG_{\mathcal{I}}\right)^{T}\right]$
$+$ $d+u^{T}_{\mathcal{A}}\left[N_{\mathcal{A}}^{T}G^{T}QAG_{\mathcal{I}}-2G_{\mathcal{A}}^{T}QGYG^{T}QG_{\mathcal{I}}+G_{\mathcal{A}}^{T}A^{T}QGN_{\mathcal{I}}\right].$
The results in this section can be directly applied to the class of index-one descriptor systems as in the nonlinear case. One obtains an explicit formula for $x_{2}$ as
$x_{2}=-[(J_{22}-R_{22})Q_{22}]^{-1}[(J_{22}-R_{22})Q_{21}+(-J_{12}^{T}-R_{12}^{T})Q_{11}]x_{1}.$
Inserting this into the first equation gives an ordinary pH system only in $x_{1}$ of the form
$\dot{x}_{1}$ $=$ $(\tilde{J}_{11}-\tilde{R}_{11})Q_{11}x_{1}+\tilde{G}_{1}u,$
$y$ $=$ $\tilde{G}_{1}^{T}Q_{11}x_{1},$
and $x_{2}$ can be determined in a post-processing step.
The cost functional is then
[…]
and thus we have again reduced the problem to one for which the previous formulas apply.
4 Conclusions and future work
For the singular optimal control problem of controlling a port-Hamiltonian system with the supplied energy as cost function, the Goh optimality conditions are derived. It is shown that the use of the specific port-Hamiltonian structure leads to very elegant solution formulas, in particular in the case of linear port-Hamiltonian systems with quadratic Hamiltonian. Optimality conditions are also presented for the case of control constraints.
It remains an open problem how to give necessary and sufficient conditions for the existence of a feedback control in the case when the matrix $W$ in the generalized Legende-Clebsch condition is only semidefinite. This is closely related to the topic to derive the optimality conditions for general port-Hamiltonian descriptor systems.
Another classical application of optimal control is the computation of (feedback) controls $u=k(x)$ or output feedbacks $u=k(y)$ so that the closed loop system is asymptotically stable. To achieve this, one just chooses an infinite interval, i.e. $T=\infty$, and requires that for the closed loop solution $\lim_{t\to\infty}x(t)=0$. For this case, analogous results to the ones obtained in this work are expected to hold. It is also an interesting question what happens when the usual approach, of adding a quadratic penalty term $\alpha\sum_{j=1}^{m}u_{i}^{2}$ to the cost function is used, and one considers the convergence behavior for $\alpha\to 0$, see [15]. The Goh conditions will not appear in that context, since $\mathcal{H}_{u}$ depends on the control and, consequently, it is, in general, not differentiable in time.
We have studied problems without terminal constraints, which admit a unique multiplier that verifies normality, i.e., the multiplier associated to the integral cost can be normalized to $1$. In [3, 4] a framework for dealing with general terminal constraints of the form $\varphi_{i}(x(T))\leq 0,$ $\eta_{j}(x(T))=0,$ for $i=1,\dots,d_{\varphi},$ and $j=1,\dots,d_{\eta}$ was introduced. Also, [12] consider terminal constraints of the form $x(T)\in K$ under an additional hypothesis.
Using the techniques in [3, 4] one may as well discuss the derivation of second order sufficient conditions.
Acknowledgment
M.S.A. acknowledges the financial support of the Brazilian agencies FAPERJ processes E-26/203.223/2017, E-26/201.346/2021 and 210.037/2024, CNPq process 310452/2019-8 and CAPES process 88881.162133/2017-01. The current research was mainly developed while M.S.A. was holding a Humboldt Fellowship at TU Berlin (Germany). V.M. has been supported by Deutsche Forschungsgemeinschaft (DFG) through the SPP1984 “Hybrid and multimodal energy systems” Project: Distributed Dynamic Security Control and by the DFG Research Center Math+, Project No. 390685689. Advanced Modeling, Simulation, and Optimization of Large Scale Multi-Energy Systems.
Appendix A Appendix
In this Appendix we recall some of the general results for singular optimal control (1.7)-(1.9) given in Section 1. We further introduce the auxiliary notation
$f(x)\coloneqq f_{0}(x)+\sum_{i=1}^{m}f_{i}(x)u_{i},$
that simplifies the expressions in several upcoming calculations.
A.1 Optimality conditions and feedback formula
In this subsection we recall the optimality conditions for problem (1.7)-(1.9), with the assumption that the control $u$ takes values in an open set $\mathbb{U}\subseteq\mathbb{R}^{m}$.
The pre-Hamilton function (or un-maximized Hamilton function) for (1.7)–(1.9) reads (leaving off the arguments)
$\mathcal{H}=p^{T}f,$ (A.1)
where the adjoint vector function $p$ maps into $\mathbb{R}^{n}$, so that the partial derivative w.r.t $u_{i}$, for $i=1,\ldots,m$, is given by
$\mathcal{H}_{u_{i}}=\partial_{u_{i}}\mathcal{H}=p^{T}f_{i}.$ (A.2)
The adjoint equation in this case is given by
$-\dot{p}^{T}=\partial_{x}\mathcal{H}=p^{T}\partial_{x}f,\quad p(T)=\nabla\Psi(x(T)).$ (A.3)
By applying the Pontryagin Maximum Principle [34], it follows that the expression (A.2) vanishes along optimal trajectories, i.e.
$\mathcal{H}_{u_{i}}=0,\ i=1,\ldots,m.$ (A.4)
Since (A.4) does not depend explicitly on the control variable, one can differentiate with respect to time and obtains
$0=\frac{d}{dt}\mathcal{H}_{u_{i}}=\dot{p}^{T}f_{i}+p^{\top}f_{i}^{\prime}\dot{x}=p^{T}\big{(}[f_{0},f_{i}]+\sum_{j=1}^{m}u_{j}[f_{j},f_{i}]\big{)},$ (A.5)
where we used the state equation (1.8) and the adjoint equation (A.3) to replace $\dot{x}$ and $\dot{p},$ and where $f_{i}^{\prime}$ denotes the Jacobian, i.e., the derivative of $f_{i}$ with respect to $x$.
The following necessary condition is well-known, see e.g. [4, 12, 40].
Theorem A.1 (Goh condition).
For the optimal control problem (1.8), the Goh condition
$p^{T}[f_{i},f_{j}]=0,\quad\text{for }i,j=1,\dots,m$ (A.6)
holds along optimal trajectories.
Applying the Goh condition (A.6) to the expression $\frac{d}{dt}\mathcal{H}_{u_{i}}$ in (A.5) yields
$\frac{d}{dt}\mathcal{H}_{u_{i}}=p[f_{0},f_{i}],$ (A.7)
so that $\frac{d}{dt}\mathcal{H}_{u_{i}}$ is differentiable with respect to $t$. Differentiating (A.7) once more, one gets
$\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j=1}^{m}u_{j}p^{T}[f_{j},[f_{0},f_{i}]].$
Thus, from $\frac{d^{2}}{dt^{2}}\mathcal{H}_{u_{i}}=0$ one obtains a linear system for $u$ given by (2.7), with the Hessian matrix $W=[w_{ij}]$ and the right hand side vector $d$, where for $i,j=1,\dots,m$,
$w_{ij}\coloneqq p[f_{j},[f_{0},f_{i}]],\qquad d_{i}\coloneqq p[f_{0},[f_{0},f_{i}]].$ (A.8)
This leads to a generalized Legendre-Clebsch condition, see e.g. [18, Theorem 6.2 and/or Corollary 6.3], [4], or [12].
Theorem A.2 (Generalized Legendre-Clebsch condition).
For the nonlinear control-affine problem (1.7)-(1.9), the associated Hessian matrix $W$ in (A.8) is negative semidefinite along optimal trajectories.
Observe that if $W$ is actually negative definite, then, from (2.7), one can express the optimal control in feedback form as
$u=-W^{-1}d.$ (A.9)
As already discussed in the introduction, the traditional technique of adding a regularizing term of the form $\alpha\sum_{i=1}^{m}u_{i}^{2}$ with positive $\alpha$ and then taking $\alpha\to 0$ does not work in this context, since (A.4) will depend on $u$ and then the expression (A.5) would not be possible. On the other hand, if one aims at obtaining a feedback formula for the controls on singular arcs by applying Pontryagin’s Maximum Principle to the modified quadratic cost and taking $\alpha\to 0,$ the expression contains $\alpha$ in the denominator and the limit may fail to exist, see, e.g. [15]. Despite this difficulty, it has been shown in the case of linear constant coefficients that this problem has some hidden regularity in a subspace [10].
Having reviewed the Goh conditions for the general unconstrained control-affine problem (1.7)-(1.9), in the next subsection we recall the optimality conditions for optimal control problems with control bounds.
A.2 Control-affine problems with control bounds
In this appendix we study problem (1.7)–(1.9) in the presence of bounds on the control as in (2.8).
In this case, the Goh optimality condition is restricted to the components that lie in the interior of the control set and one has the following optimality condition, see [4, 12].
Theorem A.3 (Goh condition under control constraints).
Consider problem (1.7)–(1.9) under the control bounds $u(t)\in\mathbb{U}$, for a.e. $t\in[0,T],$ with the set $\mathbb{U}$ as in (2.8). Then, on any open time interval over which the optimal control $u^{*}$ satisfies
$\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i}\quad\text{ and }\quad\underline{u}_{j}<u^{*}_{j}(t)<\overline{u}_{j},$
it holds that (omitting arguments)
$p^{T}\,[f_{i},f_{j}]=0.$ (A.10)
One gets the following generalized Legendre-Clebsch condition for the square submatrix of $W$ corresponding to singular components.
Theorem A.4 (Generalized Legendre-Clebsch condition).
Under the hypotheses of Theorem A.3, let $(a,b)$ be an open interval on which
$\left\{\begin{array}[]{ll}\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i},&\text{for }i\in S,\\ u^{*}_{i}(t)=\underline{u}_{i},&\text{for }i\in\underline{B},\\ u^{*}_{i}(t)=\overline{u}_{i},&\text{for }i\in\overline{B}.\end{array}\right.$
Here $\{S,\underline{B},\overline{B}\}$ is a partition of $\{1,\dots,m\}$ into singular, lower-bang and upper-bang components of the control $u^{*}.$ Then, on $(a,b)$ and for $i\in S,$ it holds that
$p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j\in S}u_{j}^{*}p^{T}[f_{j},[f_{0},f_{i}]]$ (A.11)
$+\sum_{j\in\underline{B}}\underline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]+\sum_{j\in\overline{B}}\overline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]=0.$
Additionally, the $|S|\times|S|$-matrix $W_{S}(t)$ with entries $w_{ij}=p^{T}[f_{j},[f_{0},f_{i}]],$ for $i,j\in S$, is negative semidefinite.
Proof.
The expression (A.11) follows from calculations analogous to (A.2)–(A.5), (A.7)–(2.7) and Goh condition (A.10). The negative semidefiniteness of $W_{S}$ holds in view of [12, Theorem 4.1]. ∎
Remark A.1 (A feedback formula for the singular controls).
The matrix $W_{S}$ in Theorem A.4 is the submatrix of $W$ in (A.8) corresponding to the singular components of the control. Whenever $W_{S}$ is negative definite, by using expression (2.4), one can write the singular control in terms of the state variable $x$, the adjoint variable $p$ and the bang controls.